The length of a rectangle is 8 inches more than twice its widths. If the perimeter of the rectangle is 64 in, find the width of the rectangle.

Answer by checkley71(8403)

L=2W+8
2(2W+8)+2W=64
4W+16+2W=64
6W=64-16
6W=48
W=48/6
W=8 ANSWER.
L=2*8+8
L=16+8
L=24.
PROOF
2*24+2*8=64
48+16=64
64=64